OpenStudy (anonymous):

Determine the solution set of (x + 1)2 = 25. {-8, 6} {-8, 4} {-6, 4}

OpenStudy (aaronq):

\((x + 1)^2 = 25\) \(\sqrt{(x + 1)^2} = \sqrt{25}\) \(x + 1 = \pm\sqrt{25}\) \(x + 1 = \pm 5\) can you solve it from here?

OpenStudy (anonymous):

Yes, thank you very much :)

OpenStudy (aaronq):

no problem !